# Compatible Relative Lefschetz Fibrations on Admissible Relative Stein Pairs

### v2

@@ 5. Examples
Let $D(T^{\ast}S^{n})$ denote the standard closed unit disk bundle over the $n$-sphere. The pair $(D(T^{\ast}S^{3}),D(T^{\ast}S^{2}))\subset(\mathbb{C}^{4},\mathbb{C}^{3})$ equipped with their standard complex structures (inherited from complex spaces) is an admissible relative Stein pair: It can be obtained by attaching a relative Stein $3$-handle to the trivial Stein pair $(D^{6},D^{4})$ along a relative Legendrian $2$-unknot $(L,K)\subset(S^{5},S^{3})$ embedded in the simplest possible way. So, it admits a relative Stein handle decomposition with no relative $1$- and $2$-handles, and clearly we may assume $(L,K)$ to be suspended. Therefore, following the steps in Proposition 4.2 (we take $p=q=2$), we construct a compatible relative Lefschetz fibration on the admissible relative Stein pair $(D(T^{\ast}S^{3}),D(T^{\ast}S^{2}))$ with a relative (regular) fiber $(V_{1}(2,2,2),V_{1}(2,2))\cong(D(T^{\ast}S^{2}),D(T^{\ast}S^{1}))$ and the monodromy
@@
Indeed, one can generalize the above construction: If we consider $k$-many linear plumbings of the pair $(D(T^{\ast}S^{3}),D(T^{\ast}S^{2}))$, we get the pair $(A_{k}^{6},A_{k}^{4})\cong(V_{1}(k+1,2,2),V_{1}(k+1,2))$, which is also an admissible relative Stein pair (with their standard complex structures inherited from complex spaces in which they live). It admits a relative Stein diagram as in Figure 8 where a relative Stein $3$-handle is attached along the suspended $2$-unknot $(L_{i},K_{i})$ for each $i=1,...,k$. (No relative $1$- and $2$-handles again.) Then one easily construct a compatible relative Lefschetz fibration on the admissible relative Stein pair $(A_{k}^{6},A_{k}^{4})$ with the (relative) fiber $(V_{1}(2k,2k,2),V_{1}(2k,2k))$ and the monodromy $(\widetilde{\varphi}_{2k,2k}\cdot\tau_{L_{1}}\cdots\tau_{L_{k}},\varphi_{2k,2k}\cdot\tau_{K_{1}}\cdots\tau_{K_{k}})$ where $\tau_{L_{i}},\tau_{K_{i}}$ are the right-handed Dehn twists along the zero sections of the pair $(D(T^{\ast}S^{3}),D(T^{\ast}S^{2}))$ plumbed in the $i$th place.
@@
The admissible relative Stein pair $(W,X)$ is obtained by attaching relative Stein $3$-handle to $(\widetilde{W},\widetilde{X})$ along $(\widetilde{L},\widetilde{K})$ over which the compatible relative Lefschetz fibration structure on $(\widetilde{W},\widetilde{X})$ extends. More precisely, $(W,X)$ also admits a compatible relative Lefschetz fibration with the fibers twice “punctured” Brieskorn varieties $(V_{\epsilon}(p,q,2),V_{\epsilon}(p,q))$ and with the monodromy
@@
Acknowledgments. The authors would like to thank Georgios Dimitroglou Rizell for helpful conversations and the comments on the first version.

### v3

@@ 5. Examples
Let $D(T^{\ast}S^{n})$ denote the standard closed unit disk bundle over the $n$-sphere. The pair $(D(T^{\ast}S^{3}),D(T^{\ast}S^{2}))\subset(\mathbb{C}^{4},\mathbb{C}^{3})$ equipped with their standard complex structures (inherited from complex spaces) is an admissible relative Stein pair: It can be obtained by attaching a relative $3$-handle to the trivial Stein pair $(D^{6},D^{4})$ along a relative Legendrian $2$-unknot $(L,K)\subset(S^{5},S^{3})$ embedded in the simplest possible way. So, it admits a relative Weinstein handle decomposition with no relative $1$- and $2$-handles, and clearly we may assume $(L,K)$ to be suspended. Therefore, following the steps in Proposition 4.2 (we take $p=q=2$), we construct a compatible relative Lefschetz fibration on the admissible relative Stein pair $(D(T^{\ast}S^{3}),D(T^{\ast}S^{2}))$ with a relative (regular) fiber $(V_{1}(2,2,2),V_{1}(2,2))\cong(D(T^{\ast}S^{2}),D(T^{\ast}S^{1}))$ and the monodromy
@@
Indeed, one can generalize the above construction: If we consider $k$-many linear plumbings of the pair $(D(T^{\ast}S^{3}),D(T^{\ast}S^{2}))$, we get the pair $(A_{k}^{6},A_{k}^{4})\cong(V_{1}(k+1,2,2),V_{1}(k+1,2))$, which is also an admissible relative Stein pair (with their standard complex structures inherited from complex spaces in which they live). It admits a relative Stein diagram as in Figure 7 where a relative $3$-handle is attached along the suspended $2$-unknot $(L_{i},K_{i})$ for each $i=1,...,k$. (No relative $1$- and $2$-handles again.) Then one easily construct a compatible relative Lefschetz fibration on the admissible relative Stein pair $(A_{k}^{6},A_{k}^{4})$ with the (relative) fiber $(V_{1}(2k,2k,2),V_{1}(2k,2k))$ and the monodromy $(\widetilde{\varphi}_{2k,2k}\cdot\tau_{L_{1}}\cdots\tau_{L_{k}},\varphi_{2k,2k}\cdot\tau_{K_{1}}\cdots\tau_{K_{k}})$ where $\tau_{L_{i}},\tau_{K_{i}}$ are the right-handed Dehn twists along the zero sections of the pair $(D(T^{\ast}S^{3}),D(T^{\ast}S^{2}))$ plumbed in the $i$th place.
@@
The admissible relative Stein pair $(W,X)$ is obtained by attaching relative $3$-handle to $(\widetilde{W},\widetilde{X})$ along $(\widetilde{L},\widetilde{K})$ over which the compatible relative Lefschetz fibration structure on $(\widetilde{W},\widetilde{X})$ extends. More precisely, $(W,X)$ also admits a compatible relative Lefschetz fibration with the fibers twice “punctured” Brieskorn varieties $(V_{\epsilon}(p,q,2),V_{\epsilon}(p,q))$ and with the monodromy
@@
Acknowledgments. The authors would like to thank Georgios Dimitroglou Rizell for helpful conversations and comments on the first version.